OpenStudy (anonymous):

solve the system using the subtitution method 2x+y=4 3x-5y=6 and can you explain how you got it ;)

OpenStudy (anonymous):

y = -2x + 4 3x - 5(-2x + 4) = 6 3x + 10x - 20 = 6 13x = 26 x = 2 y = -4 + 4 y = 0 (2, 0)

OpenStudy (anonymous):

how did you get o for y